Will you please solve the system by subtraction
3x-y=-7
x+y=-9
thankyou

Answer by funmath(2933)

You can solve this by adding the two equations together and eliminating the y's then solve for x:
3x-y=-7
x+y=-9
_________
4x+0y=-16
4x=-16
4x/4=-16/4
x=-4
-------------------
Now substitute x=-4 into the second equation (it's the easier of the two) and solve for y:
-4+y=-9
+4-4+y=-9+4
y=-5
-------------------
Your solution is: (-4,-5)
-------------------------
You can check this by substituting these values into both equations and see if it checks:
3(-4)-(-5)=-7
-12+5=-7
-7=-7 I think we're right!
(-4)+(-5)=-9
-9=-9 I know we're right!!!!!